Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 623     Entry time: 25 Jun 2025, 10:27
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 25.06.2025 
General Announcement:
- Martin Doering informs on the LIMP-25: 
1. Friday the summer party takes place; please deliveries with big trucks shall be re-scheduled; 
2. Cooling water 8C ramp down on July 2nd for the SA2 area; 
3. New connection to 8C cooling water for XHS on July 3-4 in SA2.
- Gerd W. comments that HXS construction ramping up next days on 30th with Inospec works on Steel and crane activities; please be careful when passing nearby; FXE construction starts next 
week for the new control room.
- Inaki De Diego says the XTD6 access turnstiles are not working; shall be back working by 15:00 with the cable being reconnected again.
- M. Prollius says XTD6 area will be fenced and entrance to XTD7 will be restricted; barriers are working again to access the site.
- XO received many reports of the unannounced effects from the HXS activities on cabling cuts on 24.06. Follow-up will be done.

CTRL:
- Nothing to report.

DA:
- Nothing to report.

EEE:
- Nothing to report.

VAC:
- Remark to instruments: FXE and MID mono will be warm-up.

XPD:
- Nothing to report.

DET:
- Intervention at SPB/SFX AGIPD with two modules replaced.

ITDM: 
- Cooling problems.

XRO:
- Not problems.

LAS:
- All accordingly to plan.

UNSYS:
- Relocation of undulators are ongoing in SA1 and SA1: control cabinets disconnected.

SPB:
- Nothing to report.

FXE:
- Started dismantling control room: it is done. 
- Ramp down starting today.

MID:
- Preparing for shutdown.

HED:
- All fine at HED.

SCS:
- Not present.

SQS:
- Nothing to report.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00